# Granular-ball Guided Masking: Structure-aware Data Augmentation

### v1

@@ Abstract
Deep learning models have achieved remarkable success in computer vision but still rely heavily on large-scale labeled data and tend to overfit when data is limited or distributions shift. Data augmentation—particularly mask-based information dropping—can enhance robustness by forcing models to explore complementary cues; however, existing approaches often lack structural awareness and risk discarding essential semantics. We propose Granular-ball Guided Masking (GBGM), a structure-aware augmentation strategy guided by Granular-ball Computing (GBC). GBGM adaptively preserves semantically rich, structurally important regions while suppressing redundant areas through a coarse-to-fine hierarchical masking process, producing augmentations that are both representative and discriminative. Extensive experiments on multiple benchmarks demonstrate consistent improvements in classification accuracy and masked image reconstruction, confirming the effectiveness and broad applicability of the proposed method. Simple and model-agnostic, it integrates seamlessly into CNNs and Vision Transformers and provides a new paradigm for structure-aware data augmentation.
@@ I Introduction
Deep learning models—ranging from Convolutional Neural Networks (CNNs) to the more advanced Vision Transformers (ViTs)—have achieved breakthrough progress in image classification [krizhevsky2012imagenet, he2016deep, dosovitskiy2020image], object detection [ren2015faster, redmon2016you, he2017mask], and semantic segmentation [long2015fully, ronneberger2015u, chen2017deeplab]. Despite these successes, such models contain millions of parameters and depend heavily on large-scale labeled data, making them prone to overfitting [huang2017densely] and poor generalization when data is limited or the distribution shifts [hendrycks2019robustness, geirhos2019imagenet]. Data augmentation serves as a fundamental strategy to expand training distributions and improve robustness.
@@
Data augmentation methods can be grouped into spatial transformations [krizhevsky2012imagenet], color distortions [szegedy2015going], and information dropping [devries2017improved]. The first two simulate natural appearance changes, while information-dropping approaches—such as Random Erasing [zhong2020random], GridMask [chen2020gridmask], and Hide-and-Seek (HaS) [singh2018hide]—intentionally remove or transform image regions to encourage complementary feature learning and reduce over-reliance on local cues. However, these methods face a persistent trade-off: too much deletion discards critical semantics, whereas too little yields trivial gains, and their random or rigid patterns lack structural awareness, making it difficult to balance preservation with effective perturbation.
Granular-ball Computing (GBC) has recently emerged as a lightweight paradigm for structural, multi-granularity modeling, demonstrating strong capability in adaptively partitioning visual space [wang2017dgcc, Xia2023GranularballCA, cheng2024gb, xie2024gbg++, quadir2024granular, 10996538]. However, current data augmentation approaches largely lack explicit structural modeling and therefore struggle to preserve global-to-local semantic consistency. By contrast, GBC naturally provides structural awareness and multi-scale representation, offering the desirable ability to identify and retain informative regions while filtering redundant background. This observation motivates us to leverage GBC as the foundation for a principled and hierarchical augmentation mechanism.
@@
  • Structure-aware masking. Unifies granular-ball computing with saliency-driven erasure to produce masks that are both discriminative and structure-preserving.
@@ II Related Work
Several research directions underpin our study. Data augmentation, including masking-based approaches, improves model robustness by diversifying training data, while granular-ball computing provides a lightweight and structurally aware representation that inspires our method.
@@
Data augmentation. Operating solely on input data, data augmentation enhances generalization without altering network structure or introducing auxiliary losses. Basic strategies such as random flipping and cropping are standard in CNNs. Building on these, Inception-preprocessing [szegedy2015going] adds random color perturbations, AutoAugment [cubuk2019autoaugment] searches for optimal augmentation policies. Information-dropping methods (e.g., Random Erasing [zhong2020random], Hide-and-Seek (HaS) [singh2018hide], GridMask [chen2020gridmask]) further remove regions to encourage complementary cues. These techniques establish a rich foundation for improving generalization across scales and model families. Yet their masking strategies are largely random or heuristic, motivating the need for more structure-aware and adaptive approaches.
@@
Masking-based augmentations. Within data augmentation, a prominent family explicitly removes or masks local image regions. Among them, Random Erasing [zhong2020random] stochastically removes rectangular areas, AutoAugment [cubuk2019autoaugment] discovers effective masking and transformation policies, GridMask [chen2020gridmask] imposes regular grid occlusions, and Hide-and-Seek (HaS) [singh2018hide] randomly hides patches at multiple scales to diversify learned features. These approaches are computationally lightweight but rely on random or heuristic patterns, which may over-delete critical semantics or provide limited regularization. They typically yield strong gains on small-scale benchmarks like CIFAR-10/100 but bring only modest improvements on large-scale datasets where structural cues dominate. In contrast, our GBGM introduces a structure-aware masking mechanism guided by granular-ball representations, bridging the gap between block-level occlusion and information-driven regularization for more stable and stronger gains.
@@
Granular-ball Computing. Human cognition follows the principle of “global precedence” [chen1982topological]. Based on traditional granular computing, Wang et al. [wang2017dgcc] proposed multi-granular cognitive computing, and Xia et al. [Xia2023GranularballCA] further introduced granular-ball computing (GBC), which represents and covers sample space with balls of varying sizes for multi-scale learning. Granular-ball computing has been applied across various artificial intelligence domains, examples include granular-ball reinforcement learning [liu2024unlock], granular-ball classifiers [xia2022efficient, xia2025graph], granular-ball clustering [xia2020ball, xie2024mgnr, xie2024gbg++], granular-ball support vector machines [quadir2024granular], and granular-ball graph representation[10996538]. Building on this concept, we develop GBGM, which leverages multi-scale granular-ball representations to guide hierarchical mask generation with strong structural awareness and low computational cost.
@@ III Method
Given an input image of size $H\times W$, GBGM progressively filters regions by local purity, propagating only the most informative areas and yielding a final binary mask $\mathcal{M}_{\text{final}}$ of size $H\times W$. At its core, granular-balls—compact structural units rather than individual pixels—capture local properties and provide the purity metric guiding selection.
Figure 2: Overview of the proposed GBGM pipeline. Granular-ball analysis guides a hierarchical selection process to retain informative regions, producing a high-resolution binary importance mask.
Formally, given a dataset $\mathcal{D}$, a granular-ball is denoted as $\mathcal{GB}_{j}$ $(j=1,2,\ldots,k)$, where $k$ is the total number of granular-balls and $\mathcal{GB}_{j}=\{\,y_{i}\mid 1\leq i\leq m_{j}\,\}$ contains $m_{j}$ data points. The center of $\mathcal{GB}_{j}$ is defined as $c_{j}=\tfrac{1}{m_{j}}\sum_{i=1}^{m_{j}}y_{i}$, and its radius as $r_{j}=\tfrac{1}{m_{j}}\sum_{i=1}^{m_{j}}\lvert y_{i}-c_{j}\rvert$. When the quality of a granular-ball fails to meet the threshold $T$, it is recursively split into smaller granular-balls until each satisfies $T$. This iterative refinement ensures convergence and yields clearer decision boundaries. In GBGM, these granular-balls serve as the structural primitives for hierarchical mask generation, as illustrated in Fig. 2.
@@ III-D Time Complexity Analysis
Algorithm 1 Granular-ball Guided Masking Pipeline
@@ IV-A1 CIFAR-10 with ResNet
We evaluate GBGM on CIFAR-10 with six ResNet [he2016deep] backbones (20/32/44/56/110/1202). Given the small image size ($32\times 32$), we adopt a single-stage mask: each image is split into $4\times 4$ patches to form an $8\times 8$ grid (Fig. 3(a)). Two masking ratios, mask10 and mask20, remove the lowest 10% and 20% of patches, with a random_mask10 baseline for comparison. Training follows standard CIFAR-10 settings: 200 epochs, batch size 128, SGD with a cosine-annealing schedule, and common augmentations (random crop, horizontal flip, normalization).
@@
Random Erasing [zhong2020random] 92.65
AutoAugment [cubuk2019autoaugment] 91.52
GridMask [chen2020gridmask] 92.68
HAS [singh2018hide] 94.35
@@
Figure 3: (a) CIFAR-10 example and (b) CIFAR-100 example. The first row shows the original input image; the second row presents the partitioning result using the granular-ball representation; the third row illustrates the saliency heatmap guided by granular balls; and the fourth row shows the final importance mask for image masking.
@@ IV-A2 CIFAR-100 with EfficientNet-L2
To further assess the generalization of GBGM, we conduct experiments on the CIFAR-100 dataset using EfficientNet-L2 [2021Sharpness] as the backbone. Because CIFAR images are small ($32\times 32$), all inputs are first resized to $224\times 224$ to match the model input resolution. We train EfficientNet-L2 for 30 epochs on a single GPU using the SAM optimizer with cosine-annealing learning rate scheduling and a weight decay of $1\times 10^{-5}$, while keeping all hyper-parameters identical across experiments. Standard data preprocessing includes resizing, random cropping, and horizontal flipping. Each baseline augmentation (AutoAugment, Random Erasing, GridMask, HAS) is tested independently for fair comparison. GBGM (Ours) is applied in a single stage similar to the CIFAR-10 experiments, as show in Fig. 3(b). Each image is divided into non-overlapping $4\times 4$ patches, and the 10% least informative patches are masked according to our importance estimation before resizing to $224\times 224$. All other training settings remain unchanged.
@@
Random Erasing [zhong2020random] 87.00 97.01
AutoAugment [cubuk2019autoaugment] 92.95 99.17
GridMask [chen2020gridmask] 85.61 96.48
HAS [singh2018hide] 85.71 96.94
@@
Figure 4: Visualization of GBGM on ImageNet-1K. From top to bottom: original images, granular-ball partitioning, saliency heatmaps, and final importance masks. The generated masks highlight semantically informative regions and suppress background, helping ViT focus on structurally important areas and improving classification accuracy.
@@ IV-A3 ImageNet-1K on CNN and Transformer Backbones
We further evaluate GBGM on the ImageNet-1K dataset with two representative architectures: a convolutional network (EfficientNet-B0) and a hierarchical vision transformer (Swin Transformer-Tiny, Swin-T) [liu2021swin], both implemented via the timm library. For EfficientNet-B0, we train for 30 epochs on a single NVIDIA GPU with batch size 64, using AdamW (learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-5}$) and a cosine-annealing learning-rate schedule. For Swin-T, we fine-tune the pretrained model on a single GPU with AdamW and a warmup+cosine-annealing schedule. In all cases, the input resolution is fixed to $224\times 224$, and the data pipeline follows standard ImageNet preprocessing: resize to $224\times 224$, random horizontal flip, mild color jitter/small rotation, and standard ImageNet normalization. For each backbone, we vary only the data augmentation strategy (AutoAugment, Random Erasing, GridMask, HAS), keeping all other hyper-parameters identical.
@@
Random Erasing [zhong2020random] 74.58 92.41 80.33 95.12
AutoAugment [cubuk2019autoaugment] 73.34 91.64 80.33 95.14
GridMask [chen2020gridmask] 73.34 91.64 80.36 95.18
HAS [singh2018hide] 74.27 92.17 80.35 95.14
@@ IV-B Image Reconstruction
To further evaluate GBGM for large-scale pretraining, we integrate it into the Masked Autoencoder (MAE) [he2022masked] with a ViT-Base backbone and train on ImageNet-100 (100 random classes from ImageNet-1K, resized to $224\times 224$). Our hierarchical masking uses patch sizes of $32\times 32$ and $16\times 16$.
@@
Method PSNR SSIM LPIPS
MAE [he2022masked] 8.22 dB 0.231 0.640
@@
IV-C Empirical Complexity Analysis

### v2

# Granular Ball Guided Masking: Structure-aware Data Augmentation

@@ Abstract
Deep learning models have achieved remarkable success in computer vision but still rely heavily on large-scale labeled data and tend to overfit when data is limited or distributions shift. Data augmentation—particularly mask-based information dropping—can enhance robustness by forcing models to explore complementary cues; however, existing approaches often lack structural awareness and risk discarding essential semantics. We propose Granular Ball Guided Masking (GBGM), a structure-aware augmentation strategy guided by Granular Ball Computing (GBC). GBGM adaptively preserves semantically rich, structurally important regions while suppressing redundant areas through a coarse-to-fine hierarchical masking process, producing augmentations that are both representative and discriminative. Extensive experiments on multiple benchmarks demonstrate consistent improvements not only in image classification and masked image reconstruction, but also in image tampering detection, validating the effectiveness and generalization of GBGM across both recognition and forensic scenarios. Simple and model-agnostic, GBGM integrates seamlessly into CNNs and Vision Transformers, offering a practical paradigm for structure-aware data augmentation.
@@ I Introduction
Deep learning models—ranging from Convolutional Neural Networks (CNNs) to the more advanced Vision Transformers (ViTs)—have achieved breakthrough progress in image classification [16, 13, 8], object detection [22, 21, 12], and semantic segmentation [19, 23, 1]. Despite these successes, such models contain millions of parameters and depend heavily on large-scale labeled data, making them prone to overfitting [15] and poor generalization when data is limited or the distribution shifts [14, 10]. Data augmentation serves as a fundamental strategy to expand training distributions and improve robustness.
@@
Data augmentation methods can be grouped into spatial transformations [16], color distortions [25], and information dropping [7]. The first two simulate natural appearance changes, while information-dropping approaches—such as Random Erasing [34], GridMask [3], and Hide-and-Seek (HaS) [24]—intentionally remove or transform image regions to encourage complementary feature learning and reduce over-reliance on local cues. However, these methods face a persistent trade-off: too much deletion discards critical semantics, whereas too little yields trivial gains, and their random or rigid patterns lack structural awareness, making it difficult to balance preservation with effective perturbation.
Granular Ball Computing (GBC) has recently emerged as a lightweight paradigm for structural, multi-granularity modeling, demonstrating strong capability in adaptively partitioning visual space [26, 31, 4, 33, 20, 6]. However, current data augmentation approaches largely lack explicit structural modeling and therefore struggle to preserve global-to-local semantic consistency. By contrast, GBC naturally provides structural awareness and multi-scale representation, offering the desirable ability to identify and retain informative regions while filtering redundant background. This observation motivates us to leverage GBC as the foundation for a principled and hierarchical augmentation mechanism.
@@
  • Structure-aware masking. Unifies Granular Ball Computing with saliency-driven erasure to produce masks that are both discriminative and structure-preserving.
@@ II Related Work
Several research directions underpin our study. Data augmentation, including masking-based approaches, improves model robustness by diversifying training data, while Granular Ball Computing provides a lightweight and structurally aware representation that inspires our method.
@@
Data augmentation. Operating solely on input data, data augmentation enhances generalization without altering network structure or introducing auxiliary losses. Basic strategies such as random flipping and cropping are standard in CNNs. Building on these, Inception-preprocessing [25] adds random color perturbations, AutoAugment [5] searches for optimal augmentation policies. Information-dropping methods (e.g., Random Erasing [34], Hide-and-Seek (HaS) [24], GridMask [3]) further remove regions to encourage complementary cues. These techniques establish a rich foundation for improving generalization across scales and model families. Yet their masking strategies are largely random or heuristic, motivating the need for more structure-aware and adaptive approaches.
@@
Masking-based augmentations. Within data augmentation, a prominent family explicitly removes or masks local image regions. Among them, Random Erasing [34] stochastically removes rectangular areas, AutoAugment [5] discovers effective masking and transformation policies, GridMask [3] imposes regular grid occlusions, and Hide-and-Seek (HaS) [24] randomly hides patches at multiple scales to diversify learned features. These approaches are computationally lightweight but rely on random or heuristic patterns, which may over-delete critical semantics or provide limited regularization. They typically yield strong gains on small-scale benchmarks like CIFAR-10/100 but bring only modest improvements on large-scale datasets where structural cues dominate. In contrast, our GBGM introduces a structure-aware masking mechanism guided by Granular Ball representations, bridging the gap between block-level occlusion and information-driven regularization for more stable and stronger gains.
@@
Granular Ball Computing. Human cognition follows the principle of “global precedence” [2]. Based on traditional granular computing, Wang et al. [26] proposed multi-granular cognitive computing, and Xia et al. [31] further introduced Granular Ball Computing (GBC), which represents and covers sample space with balls of varying sizes for multi-scale learning. Granular Ball Computing has been applied across various artificial intelligence domains, examples include Granular Ball reinforcement learning [17], Granular Ball classifiers [28, 29], Granular Ball clustering [30, 32, 33], Granular Ball support vector machines [20], and Granular Ball graph representation[6]. Building on this concept, we develop GBGM, which leverages multi-scale Granular Ball representations to guide hierarchical mask generation with strong structural awareness and low computational cost.
@@ III Method
Given an input image of size $H\times W$, GBGM progressively filters regions by local purity, propagating only the most informative areas and yielding a final binary mask $\mathcal{M}_{\text{final}}$ of size $H\times W$. At its core, Granular balls—compact structural units rather than individual pixels—capture local properties and provide the purity metric guiding selection.
Figure 2: Overview of the proposed GBGM pipeline. Granular Ball analysis guides a hierarchical selection process to retain informative regions, producing a high-resolution binary importance mask.
Formally, given a dataset $\mathcal{D}$, a Granular Ball is denoted as $\mathcal{GB}_{j}$ $(j=1,2,\ldots,k)$, where $k$ is the total number of Granular balls and $\mathcal{GB}_{j}=\{\,y_{i}\mid 1\leq i\leq m_{j}\,\}$ contains $m_{j}$ data points. The center of $\mathcal{GB}_{j}$ is defined as $c_{j}=\tfrac{1}{m_{j}}\sum_{i=1}^{m_{j}}y_{i}$, and its radius as $r_{j}=\tfrac{1}{m_{j}}\sum_{i=1}^{m_{j}}\lvert y_{i}-c_{j}\rvert$. When the quality of a Granular Ball fails to meet the threshold $T$, it is recursively split into smaller Granular balls until each satisfies $T$. This iterative refinement ensures convergence and yields clearer decision boundaries. In GBGM, these Granular balls serve as the structural primitives for hierarchical mask generation, as illustrated in Fig. 2.
@@ III-D Time Complexity Analysis
Algorithm 1 Granular Ball Guided Masking Pipeline
@@ IV-A1 CIFAR-10 with ResNet
We evaluate GBGM on CIFAR-10 with six ResNet [13] backbones (20/32/44/56/110/1202). Given the small image size ($32\times 32$), we adopt a single-stage mask: each image is split into $4\times 4$ patches to form an $8\times 8$ grid (Fig. 3(a)). Two masking ratios, mask10 and mask20, remove the lowest 10% and 20% of patches, with a random_mask10 baseline for comparison. Training follows standard CIFAR-10 settings: 200 epochs, batch size 128, SGD with a cosine-annealing schedule, and common augmentations (random crop, horizontal flip, normalization).
@@
Random Erasing [34] 92.65
AutoAugment [5] 91.52
GridMask [3] 92.68
HAS [24] 94.35
@@
Figure 3: (a) CIFAR-10 example and (b) CIFAR-100 example. The first row shows the original input image; the second row presents the partitioning result using the Granular Ball representation; the third row illustrates the saliency heatmap guided by granular balls; and the fourth row shows the final importance mask for image masking.
@@ IV-A2 CIFAR-100 with EfficientNet-L2
To further assess the generalization of GBGM, we conduct experiments on the CIFAR-100 dataset using EfficientNet-L2 [9] as the backbone. Because CIFAR images are small ($32\times 32$), all inputs are first resized to $224\times 224$ to match the model input resolution. We train EfficientNet-L2 for 30 epochs on a single GPU using the SAM optimizer with cosine-annealing learning rate scheduling and a weight decay of $1\times 10^{-5}$, while keeping all hyper-parameters identical across experiments. Standard data preprocessing includes resizing, random cropping, and horizontal flipping. Each baseline augmentation (AutoAugment, Random Erasing, GridMask, HAS) is tested independently for fair comparison. GBGM (Ours) is applied in a single stage similar to the CIFAR-10 experiments, as show in Fig. 3(b). Each image is divided into non-overlapping $4\times 4$ patches, and the 10% least informative patches are masked according to our importance estimation before resizing to $224\times 224$. All other training settings remain unchanged.
@@
Random Erasing [34] 87.00 97.01
AutoAugment [5] 92.95 99.17
GridMask [3] 85.61 96.48
HAS [24] 85.71 96.94
@@
Figure 4: Visualization of GBGM on ImageNet-1K. From top to bottom: original images, Granular Ball partitioning, saliency heatmaps, and final importance masks. The generated masks highlight semantically informative regions and suppress background, helping ViT focus on structurally important areas and improving classification accuracy.
@@ IV-A3 ImageNet-1K on CNN and Transformer Backbones
We further evaluate GBGM on the ImageNet-1K dataset with two representative architectures: a convolutional network (EfficientNet-B0) and a hierarchical vision transformer (Swin Transformer-Tiny, Swin-T) [18], both implemented via the timm library. For EfficientNet-B0, we train for 30 epochs on a single NVIDIA GPU with batch size 64, using AdamW (learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-5}$) and a cosine-annealing learning-rate schedule. For Swin-T, we fine-tune the pretrained model on a single GPU with AdamW and a warmup+cosine-annealing schedule. In all cases, the input resolution is fixed to $224\times 224$, and the data pipeline follows standard ImageNet preprocessing: resize to $224\times 224$, random horizontal flip, mild color jitter/small rotation, and standard ImageNet normalization. For each backbone, we vary only the data augmentation strategy (AutoAugment, Random Erasing, GridMask, HAS), keeping all other hyper-parameters identical.
@@
Random Erasing [34] 74.58 92.41 80.33 95.12
AutoAugment [5] 73.34 91.64 80.33 95.14
GridMask [3] 73.34 91.64 80.36 95.18
HAS [24] 74.27 92.17 80.35 95.14
@@ IV-B Image Reconstruction
To further evaluate GBGM for large-scale pretraining, we integrate it into the Masked Autoencoder (MAE) [11] with a ViT-Base backbone and train on ImageNet-100 (100 random classes from ImageNet-1K, resized to $224\times 224$). Our hierarchical masking uses patch sizes of $32\times 32$ and $16\times 16$.
@@
Method PSNR ↑ SSIM ↑ LPIPS ↓
MAE [11] 8.22 dB 0.231 0.640
@@
IV-C Image Tampering Detection
Image tampering detection aims to determine whether an image has been manipulated and to localize the tampered regions. This task is particularly challenging when the test distribution differs from the training data due to changes in manipulation sources, post-processing operations, or imaging conditions. To validate the effectiveness of GBGM beyond classification, we follow the ADCD-based tampering detection protocol in [27] and build our detector upon the same baseline model. Specifically, we adopt ADCD as the baseline and plug in GBGM to form ADCDNet_GBGM, and we further compare with several representative augmentation-based variants, including Random Erasing [34], AutoAugment [5], GridMask [3], and HAS [24]. Table VII reports the quantitative results on three validation sets, i.e., TestSet, FCD, and SCD. Overall, ADCDNet_GBGM achieves the highest average F1 of 91.36%, improving upon the baseline (90.30%) by 1.06 points, and also outperforming Random Erasing (90.71%), AutoAugment (90.71%), GridMask (90.58%), and HAS (90.41%). On the more challenging SCD set, GBGM increases F1 from 86.27% to 87.77% (+1.50), while also improving F1 on TestSet (from 91.97% to 92.65%, +0.68) and FCD (from 92.65% to 93.66%, +1.01). Notably, HAS attains the highest recall across all three sets, whereas GBGM yields the best precision and F1, indicating that GBGM improves overall detection quality under both in-domain and cross-domain evaluation settings.
TABLE VII: Performance comparison on different validation sets (%).
Method Val Set Pre Recall F1 Avg. F1
Baseline [27] TestSet 90.16 93.86 91.97 90.30
FCD 89.47 96.07 92.65
SCD 79.93 93.71 86.27
Random Erasing [34] TestSet 89.76 95.00 92.31 90.71
FCD 91.39 95.13 93.23
SCD 79.09 95.67 86.59
AutoAugment [5] TestSet 89.76 95.00 92.31 90.71
FCD 91.39 95.13 93.23
SCD 79.09 95.67 86.59
GridMask [3] TestSet 90.13 94.29 92.17 90.58
FCD 90.48 95.55 92.95
SCD 79.83 94.66 86.62
HAS [24] TestSet 86.23 95.97 90.84 90.41
FCD 88.80 97.11 92.77
SCD 80.40 96.31 87.64
GBGM (Ours) TestSet 91.15 94.19 92.65 91.36
FCD 93.54 93.78 93.66
SCD 81.52 95.06 87.77
IV-D Empirical Complexity Analysis